# Efficient quantum Gibbs samplers with Kubo–Martin–Schwinger detailed balance condition

### v2

@@ 1.2. Contribution
Table 1. A comparison of quantum Gibbs samplers using techniques related to Lindblad dynamics. Here the truncation time $T$ is used to truncate the integral in Eq. 1.1 to $[-T,T]$ in simulation. The number of jump operators for the Libladian is denoted by $\infty$ if $\gamma(\omega)$ is a continuous function in Eq. 1.2. The total cost refers to the total Hamiltonian simulation time using the best available Lindblad simulation algorithm. The $o(1)$ factor in the truncation time of this work stems from our choice of the filtering function and $o(1)$ can be chosen to be arbitrarily small without much added cost. In this work, the factor $\beta^{2}S$ scaling in the total cost is due to that our jump operator allows the Bohr frequency difference to be of size $S$ (see 23). When $S=\mathcal{O}(1/\beta)$, this recovers the linear dependence on $\beta$ in the total cost. When $S=\mathcal{O}(1)$, the dependence on $\beta$ is quadratic. The mixing time $t_{\mathrm{mix}}$ is method dependent, and their values are generally difficult to compare with each other in theory. The weak coupling assumption may be unphysical for large quantum systems. The rounding promise-based method prepares an ensemble of density operators, and the total cost for preparing each density operator in the ensemble may be improved to $\widetilde{\mathcal{O}}(t_{\rm mix}\beta\epsilon^{-2})$ (see [RWW23, Remark of Theorem 1]).
@@ 1.3. Discussion and open questions
Although rigorous bounding of the mixing time has been achieved for certain quantum Gibbs samplers operating on commuting Hamiltonians [KB16, BCG+23], establishing the mixing time for non-commuting Hamiltonians at moderate or even low temperatures presents a substantial theoretical challenge. There are two interesting works along this line. The first is that Rouzé et al [RFA24] established the spectral gap of KMS detailed balanced Lindbladians for certain $k$-local Hamiltonians at high temperatures using Lieb-Robinson estimates. The analysis in [RFA24] may be applicable in our setting and we plan to investigate this in detail in a future work. On the other hand, Bakshi et al [BLMT24] demonstrated that at the high temperature, the Gibbs State of certain $k$-local Hamiltonian becomes a linear combination of tensor products of stabilizer states, can be prepared in polynomial time using randomized classical algorithms. This suggests that exploring the relationship between the complexity of Gibbs states and mixing times could be a fruitful avenue for future research.
It is also noteworthy that introducing a coherent term to any detailed balanced Lindbladian disrupts the detailed balance condition, but the Gibbs state remains a fixed point. The influence of the coherent term on the mixing time may be significant and its characterization remains an open question. Finally, the mixing time $t_{\rm mix}$ may be very different across different quantum Gibbs samplers. Both theoretical and numerical evidence are needed in order to quantify the mixing time and to compare the efficiency of quantum Gibbs samplers for physical systems of interest.
@@ Acknowledgement
This material is based upon work supported by the Challenge Institute for Quantum Computation (CIQC) funded by National Science Foundation (NSF) through grant number OMA-2016245 (Z.D.), National Science Foundation (NSF) award under grant number DMS-2012286 and CHE-2037263 (B.L.), the Applied Mathematics Program of the US Department of Energy (DOE) Office of Advanced Scientific Computing Research under contract number DE-AC02-05CH1123, and a Google Quantum Research Award (L.L.). L.L. is a Simons investigator in Mathematics. We thank Anthony Chen, Li Gao, Marius Junge and Jianfeng Lu for insightful discussions.
Note: In completing this work, we became aware of the concurrent research by Chen, Doriguello, and Gilyén that similarly aims to develop quantum Gibbs samplers with a finite number of jump operators.
@@ 2. Structures of detailed balanced Lindbladians
We first recall that a quantum channel $\Phi:\mathcal{B}(\mathcal{H})\to\mathcal{B}(\mathcal{H})$ is a completely positive trace preserving (CPTP) map, while a quantum Markov semigroup (QMS) $(\mathcal{P}_{t})_{t\geq 0}:\mathcal{B}(\mathcal{H})\to\mathcal{B}(\mathcal{H})$, also called Lindblad dynamics, is defined as a $C_{0}$-semigroup of completely positive, unital maps. The generator
$\mathcal{L}(X):=\lim_{t\to 0}t^{-1}(\mathcal{P}_{t}(X)-X)$
@@
with $\mathcal{J}$ being the index set with cardinality $|\mathcal{J}|\leq N^{2}$.
The operators $L_{j}$ in (2.2) are called jump operators, which are non-unique for a given Lindbladian. From $\mathcal{L}({\bf 1})=0$, by (2.1), the operator $K$ can be written as
(2.3) $K=V-iG\quad\text{with}\quad V=-\frac{1}{2}\sum_{j\in\mathcal{J}}L_{j}^{\dagger}L_{j}\,,$
where $V:=\frac{K^{\dagger}+K}{2}$ and $G:=\frac{K^{\dagger}-K}{2i}$ are self-adjoint operators, and then there holds
$\mathcal{L}(X)=i[G,X]+\sum_{j\in\mathcal{J}}\left(L_{j}^{\dagger}XL_{j}-\frac{1}{2}\left\{L_{j}^{\dagger}L_{j},X\right\}\right)\,,$
where $i[G,X]$ and $\mathcal{L}(X)-i[G,X]$ are the coherent and dissipative parts of the dynamic, respectively.
@@
equivalently, the irreducible QMS $\mathcal{P}^{\dagger}_{t}$ [Wol12, Proposition 7.5]. For the reader’s convenience, we recall the definition of irreducibility and some further equivalent conditions. We say that a quantum channel $\Phi$ is irreducible if all the orthogonal projections $P$ satisfying $\Phi(P\mathcal{B}(\mathcal{H})P)\subset P\mathcal{B}(\mathcal{H})P$ are trivial, i.e., zero or identity. The following results are adapted from [Wol12, ZB23].
@@
  • The multiplicative algebra generated by the jump operators $\{L_{j}\}$ and $K=-\frac{1}{2}\sum_{j}L_{j}^{\dagger}L_{j}-iG$ gives the whole algebra $\mathcal{B}(\mathcal{H})$.
  • The operators $\{L_{j}\}$ and $K$ have no trivial common invariant subspace.
@@
We next discuss the quantum detailed balance condition (DBC), which provides a sufficient criterion to guarantee the Lindbladian’s steady state. For a given $\sigma\in\mathcal{D}_{+}(\mathcal{H})$, we define the modular operator:
@@
We also let $L_{\sigma}(X)=\sigma X$ and $R_{\sigma}(X)=X\sigma$ be the left and right multiplication operators, respectively. Then, for any $f:(0,\infty)\to(0,\infty)$ satisfying $f(1)=1$ and $\sigma\in\mathcal{D}_{+}(\mathcal{H})$, we define the following operator:
@@ 2.1. Davies generator and GNS-detailed balance
with $\mathcal{Z}_{\beta}=\operatorname{tr}(e^{-\beta H})$ being the normalization constant, called partition function. It is easy to see that any full-rank quantum state can be written as a Gibbs state $\sigma=e^{-h}$ with $h=-\log(\sigma)$.
Recall that the main aim of this work is to develop an efficient quantum Gibbs sampler via QMS. An important class of Lindbladians for this purpose are Davies generators, which describe the weak coupling limit of a system coupled to a large thermal bath [Dav76, Dav79]. It has natural applications in thermal state preparations but with inherent difficulties from the energy-time uncertainty principle [ML20, RWW23, CKBG23]; see 7. We next review the canonical form of Davies semigroups and show that they essentially characterize the Lindbladians with GNS-DBC [KFGV77].
@@
Following [CKBG23], we introduce the weighted operator Fourier Transform, which is crucial for our following algorithmic design and its analysis. Given an operator $A\in\mathcal{B}(\mathcal{H})$ and a filter function $f:\mathbb{R}\to\mathbb{C}$ with certain regularity (e.g., $L^{1}$ integrable or tempered distribution), we define
(2.15) $\widehat{A}_{f}(\omega):=\int_{-\infty}^{\infty}A(t)e^{-i\omega t}f(t)\mathrm{d}t=2\pi\sum_{\nu\in B_{H}}A_{\nu}\widehat{f}(\omega-\nu)\,,$
where
$\widehat{f}(\omega)=\frac{1}{2\pi}\int_{-\infty}^{\infty}f(t)e^{-i\omega t}\mathrm{d}t$
is the Fourier Transform of $f(t)$. In the case of $f(t)=1$, we have $\widehat{f}(\omega)=\delta(\omega)$ and
(2.16) $\widehat{A}_{f=1}(\omega)=2\pi\sum_{\nu\in B}A_{\nu}\delta(\omega-\nu)\,.$
@@
that is, $A_{\nu}^{a}$ is an eigenvector of $\Delta_{\sigma_{\beta}}$ with the eigenvalue $e^{-\beta\nu}$. Note that the condition (2.21) holds by the definition of $A_{\nu}^{a}$, while the condition (2.20) is often referred to as KMS condition [^2] The KMS condition should not be confused with the KMS detailed balance condition. These two terms are mathematically unrelated. [KFGV77], which, as we shall see below, ensures the GNS-reversibility of the Lindbladian.
@@
with $\omega_{j}\in\mathbb{R}$ and $|\mathcal{J}|\leq N^{2}-1$, where $L_{j}\in\mathcal{B}(\mathcal{H})$ satisfies
@@
with normalization constants $c_{j}>0$, and for each $j$, there exists $j^{\prime}\in\mathcal{J}$ such that
@@
with the dissipative part exactly satisfying the conditions in 6 by re-indexing $j=(a,\nu)$.
@@
The Davies generator $\mathcal{L}_{\beta}$ in (2.17) can be viewed as a quantum analog of a classical Markov chain, and therefore becomes a natural candidate for the Gibbs state preparation [RWW23, CKBG23]. However, implementing the Davies generator accurately requires being able to resolve and distinguish between all Bohr frequencies $\nu$, while the gap between two Bohr frequencies $\nu,\nu^{\prime}$ could be exponentially small as the system size increases in a generic setting. In view of (2.15) and (2.16), by the energy-time uncertainty principle, this means an impractically long Hamiltonian simulation time and is a key obstacle in leveraging the Davies semigroup directly as a quantum Gibbs sampler.
It is also worth mentioning that the sum over Bohr frequencies (2.17) is derived from a secular approximation [BP02], which may be regarded as theoretical evidence that the GNS detailed balance is an idealized construction and thus is difficult to be exactly implemented in practice.
@@ 2.2. KMS-detailed balanced generators
Recalling that the KMS DBC is a weaker property compared to the GNS one, but can still guarantee the Gibbs state as a fixed point of the dynamic, one may expect that the KMS-detailed balanced QMS can provide a more efficient class of Gibbs state preparation algorithms. In this section, we introduce the canonical form of the QMS satisfying $\sigma_{\beta}$-KMS DBC with a proof sketch, mainly following [FU07, AC21], which are fundamental for the subsequent discussion on quantum algorithms.
@@
and denote by $\mathfrak{H}_{S}^{\perp}$ its orthogonal complement. The following useful lemma characterizes the freedom of $X,Y$ in (2.25) [^3] This result is from [AC21, Lemma 3.10 and Remark 3.11], which include some typos in the arguments. We provide a short proof here for the reader’s convenience..
@@
Let $\Phi$ be a superoperator with the representation (2.25). If some $X^{\prime},Y^{\prime}\in\mathcal{B}(\mathcal{H})$ gives the same $\Phi$, then $X^{\prime}=X+\eta{\bf 1}$ and $Y^{\prime}=Y-\eta{\bf 1}$ for some $\eta\in\mathbb{C}$.
@@
Let $F_{\alpha}$ with $\alpha=(i,j)$ and $1\leq i,j\leq N$ be a basis of $\mathcal{B}(\mathcal{H})$ satisfying $\langle F_{\alpha},F_{\beta}\rangle/N=\delta_{\alpha\beta}$, $F_{(1,1)}={\bf 1}$, and $F^{\dagger}_{(i,j)}=F_{(j,i)}$, and let $E_{i,j}$ be another basis defined by $E_{i,j}=\sqrt{N}\ket{u_{i}}\bra{u_{j}}$ |, where $\{u_{j}\}$ is an orthonormal basis of $\mathcal{H}$. Suppose that $\Phi(A)=XA+AY$ for some $X,Y$. We compute the coefficients for the expansion of $\Phi(A)=\sum_{\alpha,\beta}(C_{\Phi})_{\alpha,\beta}F_{\alpha}^{\dagger}AF_{\beta}$:
$(C_{\Phi})_{\alpha,\beta}$ $=\frac{1}{N^{3}}\sum_{i,j=1}^{N}\operatorname{tr}\left[(F_{\alpha}^{\dagger}E_{i,j}F_{\beta})^{\dagger}\Phi(E_{i,j})\right]$
@@
Then, by 8, we derive $K^{\dagger}=\Delta_{\sigma_{\beta}}^{-1/2}K+\eta{\bf 1}$ for some $\eta\in\mathbb{C}$, where $\eta$ must be zero, thanks to $\operatorname{tr}(K^{\dagger})=\operatorname{tr}(K)=\operatorname{tr}(\Delta_{\sigma_{\beta}}^{-1/2}K)\in\mathbb{R}$. The proof is complete. ∎
We proceed to derive an explicit formula for the operator $K$. Recalling the decomposition (2.3) and $V=-\frac{1}{2}\sum_{j\in\mathcal{J}}L_{j}^{\dagger}L_{j}$, it suffices to find an expression for the involved operator $G$. To do so, we reformulate the constraint (2.29) above as a Lyapunov equation:
@@
and $\tanh(\log(x^{1/4}))=\frac{x^{1/2}-1}{x^{1/2}+1}$. Then, by functional calculus and (2.30), there holds
@@
with $L_{j}$ satisfying (2.28), and $G$ being self-adjoint and given by

### v4

@@ 1.2. Contribution and outline
Table 1. A comparison of some quantum Gibbs samplers using techniques related to Lindblad dynamics with $\beta>1$. Here the truncation time $T$ is used to truncate the integral in Eq. 1.4 to $[-T,T]$ in simulation. For each jump proposal $A^{a}$, the number of jump operators for the Lindbladian is denoted by $\infty$ if $\gamma(\omega)$ is a continuous function in Eq. 1.5. The total cost refers to the total Hamiltonian simulation time using the best available Lindblad simulation algorithm. The $o(1)$ factor in the truncation time of this work stems from our choice of the filtering function and $o(1)$ can be chosen to be arbitrarily small without much added cost. The mixing time $t_{\mathrm{mix}}$ is method-dependent, and its value is generally difficult to compare with each other in theory. The weak coupling assumption may be unphysical for large quantum systems. The rounding promise-based method prepares an ensemble of density operators, and the total cost for preparing each density operator in the ensemble may be improved to $\widetilde{\mathcal{O}}(\beta t_{\rm mix}\epsilon^{-2})$ by the improved Lindbladian simulation technique from [CKBG23] (see [RWW23, Remark of Theorem 1]).
@@ 1.3. Discussion and open questions
[…]
We emphasize that in this work, an efficient quantum Gibbs sampler means that the proposed Gibbs sampler can be efficiently simulated on a quantum computer, while in classical MCMC literature, efficient often indicates rapid mixing. The rigorous bounding of the mixing time has been achieved for Davies generators for some general families of local commuting Hamiltonians [KB16, BCG+23, KACR24], whereas establishing the mixing time for non-commuting Hamiltonians at moderate or even low temperatures presents a substantial theoretical challenge. There are two interesting works along this line. The first is that Rouzé et al [RFA24] established the spectral gap of KMS detailed balanced Lindbladians for certain $k$-local Hamiltonians at high temperatures using Lieb-Robinson estimates. The analysis in [RFA24] may be applicable in our setting and we plan to investigate this in detail in a future work. On the other hand, Bakshi et al [BLMT24] demonstrated that at the high temperature, the Gibbs State of certain $k$-local Hamiltonian becomes a linear combination of tensor products of stabilizer states, can be prepared in polynomial time using randomized classical algorithms. This suggests that exploring the relationship between the complexity of Gibbs states and mixing times could be a fruitful avenue for future research.
It is also noteworthy that introducing a nontrivial coherent term $i[G,\cdot]$ with $[G,\sigma_{\beta}]=0$ to any $\sigma_{\beta}$-KMS detailed balanced Lindbladian potentially disrupts the detailed balance condition, but the Gibbs state remains a fixed point. The influence of the coherent term on the mixing time may be significant and its characterization remains a largely open question. We refer interested readers to [FLT24, LL24] for the recent progress on analyzing the coherent term effect via hypocoercivity. Finally, the mixing time $t_{\rm mix}$ may be very different across different quantum Gibbs samplers. Both theoretical and numerical evidence are needed in order to quantify the mixing time and to compare the efficiency of quantum Gibbs samplers for physical systems of interest.
@@ Acknowledgement
This material is based upon work supported by the Challenge Institute for Quantum Computation (CIQC) funded by National Science Foundation (NSF) through grant number OMA-2016245 (Z.D.), National Science Foundation (NSF) award under grant number DMS-2012286 and CHE-2037263 (B.L.), the Applied Mathematics Program of the US Department of Energy (DOE) Office of Advanced Scientific Computing Research under contract number DE-AC02-05CH1123, and a Google Quantum Research Award (L.L.). L.L. is a Simons investigator in Mathematics. We thank Anthony Chen, András Gilyén, Li Gao, Marius Junge and Jianfeng Lu for insightful discussions.
Note: In completing this work, we became aware of the concurrent research by Chen, Doriguello, and Gilyén [GCDK24], which similarly aims to develop efficiently simulable discrete-time quantum Gibbs samplers, in analog with the classical Metropolis sampling.
@@ 2. Structures of detailed balanced Lindbladians
We first recall that a quantum channel $\Phi:\mathcal{B}(\mathcal{H})\to\mathcal{B}(\mathcal{H})$ is a completely positive trace preserving (CPTP) map, and a quantum Markov semigroup (QMS)222 With some abuse of terminology, we will refer to both $\mathcal{P}_{t}$ and its adjoint dynamics $\mathcal{P}^{\dagger}_{t}$ as a QMS. $(\mathcal{P}_{t})_{t\geq 0}:\mathcal{B}(\mathcal{H})\to\mathcal{B}(\mathcal{H})$, also called Lindblad dynamics, is defined as a $C_{0}$-semigroup of completely positive, unital maps. The generator
$\mathcal{L}(X):=\lim_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}t\to 0^{+}}}t^{-1}(\mathcal{P}_{t}(X)-X)$
@@
with $\mathcal{J}$ being an index set with cardinality $|\mathcal{J}|\leq N^{2}$. In particular, letting $G:=\frac{K^{\dagger}-K}{2i}$, we have the Lindbladian form:
where $i[G,X]$ and $\mathcal{L}(X)-i[G,X]$ are the coherent and dissipative parts of dynamics, respectively.
The operators $L_{j}$ in (2.3) are called jump operators, which are non-unique for a given Lindbladian. The formula (2.3) is a direct consequence of (2.1). In fact, from $\mathcal{L}({\bf 1})=0$ with (2.1), the operator $K$ can be written as $K=V-iG$ with $V=-\frac{1}{2}\sum_{j\in\mathcal{J}}L_{j}^{\dagger}L_{j}$, where $V:=\frac{K^{\dagger}+K}{2}$ and $G:=\frac{K^{\dagger}-K}{2i}$ are self-adjoint operators. Plugging this into (2.1) readily gives (2.3).
@@
equivalently, the irreducible QMS $\mathcal{P}^{\dagger}_{t}$ [Wol12, Proposition 7.5]. For readers’ convenience, we recall the definition of irreducibility and some equivalent conditions. We say that a quantum channel $\Phi$ is irreducible if all the orthogonal projections $P$ satisfying $\Phi(P\mathcal{B}(\mathcal{H})P)\subset P\mathcal{B}(\mathcal{H})P$ are trivial, i.e., zero or identity. The following results are adapted from [Wol12, ZB23].
@@
  • The $\mathbb{C}$-algebra generated by the jump operators $\{L_{j}\}_{j\in\mathcal{J}}$ and $K=-\frac{1}{2}\sum_{j}L_{j}^{\dagger}L_{j}-iG$ gives the whole algebra $\mathcal{B}(\mathcal{H})$.
  • The operators $\{L_{j}\}_{j\in\mathcal{J}}$ and $K$ have no trivial common invariant subspace.
@@
We next discuss the quantum detailed balance condition (DBC), which provides a sufficient criterion to guarantee steady states of a Lindbladian. For a given full-rank state $\sigma\in\mathcal{D}_{+}(\mathcal{H})$, we define the modular operator:
@@
We also let $L_{\sigma}(X)=\sigma X$ and $R_{\sigma}(X)=X\sigma$ be the left and right multiplication operators, respectively. Then, for any $f:(0,\infty)\to(0,\infty)$ satisfying $f(1)=1$ and $\sigma\in\mathcal{D}_{+}(\mathcal{H})$, we define the operator:
@@ 2.1. Davies generator and GNS-detailed balance
with $\mathcal{Z}_{\beta}=\operatorname{tr}(e^{-\beta H})$ being the normalization constant called the partition function. It is easy to see that any full-rank quantum state can be written as a Gibbs state $\sigma=e^{-h}$ with $h=-\log(\sigma)$.
Recall that our main aim is to design an efficient quantum Gibbs sampler via QMS. An important class of Lindbladians for this purpose are Davies generators, which describe the weak coupling limit of a system coupled to a large thermal bath [Dav76, Dav79]. It has natural applications in thermal state preparations but with inherent difficulties from the energy-time uncertainty principle [ML20, RWW23, CKBG23]; see 7 below. We next review the canonical form of Davies generators and show that they essentially characterize the Lindbladians with GNS-DBC [KFGV77].
@@
For our later use, we introduce a weighted integral $A_{f}$ of the Heisenberg evolution $A(t)$ with respect to an $L^{1}$-integrable function $f$: for $A\in\mathcal{B}(\mathcal{H})$,
(2.15) $A_{f}:=\int_{-\infty}^{\infty}A(t)f(t)\,\mathrm{d}t=\sum_{\nu\in B_{H}}A_{\nu}\widehat{f}(\nu)\,,$
where $\widehat{f}(\nu)$ is the inverse Fourier transform of $f(t)$:
$\widehat{f}(\nu)=\int_{-\infty}^{\infty}f(t)e^{i\nu t}\,\mathrm{d}t\,.$
Moreover, by a direct computation, we have
(2.16) $A_{f}=\widehat{f}(-\beta^{-1}\log\Delta_{\sigma_{\beta}})A\,,$
where the operator $\widehat{f}(-\beta^{-1}\log\Delta_{\sigma_{\beta}})$ is defined by functional calculus based on the modular operator $\Delta_{\sigma}$, noting that the Fourier transform of an $L^{1}$ function is bounded and uniformly continuous.
@@
that is, $A_{\nu}^{a}$ is an eigenvector of $\Delta_{\sigma_{\beta}}$ with the eigenvalue $e^{-\beta\nu}$. Note that the condition (2.21) holds by the definition of $A_{\nu}^{a}$, while the other one (2.20) is often referred to as KMS condition [^3] The KMS condition should not be confused with the KMS detailed balance condition. These two terms are mathematically unrelated. [KFGV77]. These two conditions, as we shall see below, ensure the GNS-reversibility of the Lindbladian.
@@
with $\omega_{j}\in\mathbb{R}$, where $\mathcal{J}$ is an index set with cardinality $|\mathcal{J}|\leq N^{2}-1$, and $L_{j}\in\mathcal{B}(\mathcal{H})$ satisfies
@@
for normalization constants $c_{j}>0$, and for each $j$, there exists $j^{\prime}\in\mathcal{J}$ such that
@@
where the dissipative part exactly satisfies the conditions in 6 by re-indexing $j=(a,\nu)$.
@@
As reviewed in the introduction, the Davies generator $\mathcal{L}_{\beta}$ in (2.17) has been adopted for the thermal state preparation [RWW23, CKBG23, CKBG23]. However, implementing it with a small error requires accurately resolving all the Bohr frequencies $\nu$, while the minimal gap between two Bohr frequencies $\nu,\nu^{\prime}$ could be exponentially small in the worst case. By the energy-time uncertainty principle, this means an impractically long Hamiltonian simulation time and is a key obstacle in directly leveraging the Davies semigroup as an efficient quantum Gibbs sampler [CKBG23].
It is also worth mentioning that the sum over Bohr frequencies (2.17) is derived from a secular approximation [BP02], which may be regarded as theoretical evidence that the GNS detailed balance is an idealized construction and thus is difficult to be efficiently implemented with desired accuracy.
@@ 2.2. KMS-detailed balanced generators
Recalling that the KMS DBC is a weaker property compared to the GNS one, but can still guarantee the Gibbs state as a fixed point of the dynamics, one may expect that the KMS-detailed balanced QMS can provide a more efficient class of Gibbs state preparation algorithms. The first complete characterization of the $\sigma$-KMS detailed balanced Lindbladian was given in [FU07, Theorems 7.2 and 7.3], where some equivalent conditions on the coherent term and jump operators were proposed. In this section, we revisit their arguments along the lines of the recent work [AC21] and introduce the canonical form of the QMS satisfying $\sigma_{\beta}$-KMS DBC with a self-contained proof, which are the starting points of our quantum Gibbs samplers.
@@
and denote by $\mathfrak{H}_{S}^{\perp}$ its orthogonal complement for the normalized Hilbert–Schmidt inner product on superoperators, which is defined as follows: for $\Phi,\Psi\in\mathfrak{H}$,
$\langle\Phi,\Psi\rangle_{\mathfrak{H}}=\frac{1}{N^{2}}\sum_{i,j=1}^{N}\langle\Phi(X_{i,j}),\Phi(X_{i,j})\rangle\,,$
with $\{X_{i,j}\}_{i,j=1}^{N}$ being any orthonormal basis of $\mathcal{B}(\mathcal{H})$ satisfying
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\langle X_{i,j},X_{k,l}\rangle=\operatorname{tr}(X_{i,j}^{\dagger}X_{k,l})=\delta_{(i,j),(k,l)}\,.}$
The following useful lemma characterizes the freedom of $X,Y$ in (2.25) [^4] This result is from [AC21, Lemma 3.10 and Remark 3.11], which include some typos in the arguments. We provide a short proof here for the reader’s convenience..
@@
Let $\Phi$ be a superoperator with the representation (2.25). If some $X^{\prime},Y^{\prime}\in\mathcal{B}(\mathcal{H})$ gives the same $\Phi$ via (2.25), then $X^{\prime}=X+\eta{\bf 1}$ and $Y^{\prime}=Y-\eta{\bf 1}$ for some $\eta\in\mathbb{C}$.
@@
Let $F_{\alpha}$ with $\alpha=(i,j)$ and $1\leq i,j\leq N$ be a basis of $\mathcal{B}(\mathcal{H})$ satisfying $\langle F_{\alpha},F_{\beta}\rangle/N=\delta_{\alpha\beta}$, $F_{(1,1)}={\bf 1}$, and $F^{\dagger}_{(i,j)}=F_{(j,i)}$, and let $E_{i,j}$ be another basis defined by $E_{i,j}=\ket{u_{i}}\bra{u_{j}}$ |, where $\{u_{j}\}$ is an orthonormal basis of $\mathcal{H}$. Suppose that $\Phi(A)=XA+AY$ for some $X,Y$. Note that $\Phi_{\alpha,\beta}(X)=F_{\alpha}^{\dagger}AF_{\beta}$ is an orthonormal basis of $\mathfrak{H}$ by computing
$\langle\Phi_{\alpha,\beta},\Phi_{\gamma,\eta}\rangle_{\mathfrak{H}}=\frac{1}{N}\langle F_{\gamma},F_{\alpha}\rangle\frac{1}{N}\langle F_{\beta},F_{\eta}\rangle=\delta_{(\alpha,\beta),(\gamma,\eta)}\,.$
We then compute the coefficients for the orthogonal expansion of $\Phi(A)=\sum_{\alpha,\beta}(C_{\Phi})_{\alpha,\beta}F_{\alpha}^{\dagger}AF_{\beta}$:
$(C_{\Phi})_{\alpha,\beta}$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{N^{2}}}\sum_{i,j=1}^{N}\operatorname{tr}\left[(F_{\alpha}^{\dagger}E_{i,j}F_{\beta})^{\dagger}\Phi(E_{i,j})\right]$
@@
Then, by 8, we derive $K^{\dagger}=\Delta_{\sigma_{\beta}}^{-1/2}K+\eta{\bf 1}$ for some $\eta\in\mathbb{C}$, where $\eta$ must be zero, thanks to $\operatorname{tr}(K^{\dagger})=\operatorname{tr}(K)=\operatorname{tr}(\Delta_{\sigma_{\beta}}^{-1/2}K)\in\mathbb{R}$. The proof is complete.
We proceed to derive an explicit formula for the operator $K$. Recalling the decomposition $K=V-iG$ with $V=-\frac{1}{2}\sum_{j\in\mathcal{J}}L_{j}^{\dagger}L_{j}$ ( see after 1), it suffices to find an expression for $G$. To do so, we reformulate the constraint (2.29) above as a Lyapunov equation:
@@
and $\tanh(\log(x^{1/4}))=\frac{x^{1/2}-1}{x^{1/2}+1}$. Then, by functional calculus and (2.30), it holds that
@@
with $L_{j}$ satisfying
$\Delta_{\sigma_{\beta}}^{-1/2}L_{j}=L_{j}^{\dagger}\,,$
and $G$ being self-adjoint and given by